OpenStudy (ashley1nonly):

Determine the value of Kp for the following reaction if the equilibrium concentrations are as follows: P(CO)eq = 6.8 × 10-11 atm, P(O2)eq = 1.3 × 10-3 atm, P(CO2)eq = 0.041 atm. 2 CO(g) + O2(g) ⇌ 2 CO2(g) A) 3.6 × 10-21 B) 2.8 × 1020 C) 4.6 × 1011 D) 2.2 × 10-12 E) 3.6 × 10-15 Answer: B

OpenStudy (ashley1nonly):

I know how to find Kc but i cant figure out how to solve for Kp

OpenStudy (anonymous):

use product over reactant, no solids or liquids

OpenStudy (anonymous):

Il work it out 1sec, its been a while since chem 2

OpenStudy (ashley1nonly):

kp= P^2 [CO2] / P^2[ CO2] P [O2]

OpenStudy (ashley1nonly):

ok

OpenStudy (anonymous):

Perfect yes, just plug and solve

OpenStudy (ashley1nonly):

Thats it

OpenStudy (anonymous):

yep, wait till you take analytic chem there is more steps

OpenStudy (ashley1nonly):

With the P^2 would that mean square it

OpenStudy (anonymous):

yes

OpenStudy (ashley1nonly):

ok thanks very much

OpenStudy (anonymous):

np anymore questions feel free to ask

OpenStudy (ashley1nonly):

yes

OpenStudy (ashley1nonly):

when do we know the kp=Kc

OpenStudy (anonymous):

1 sec let me use the rr

OpenStudy (ashley1nonly):

ok

OpenStudy (ashley1nonly):

For the reaction: N2(g) + 2 O2(g) ⇌ 2 NO2(g), Kc = 8.3 × 10-10 at 25°C. What is the concentration of N2 gas at equilibrium when the concentration of NO2 is five times the concentration of O2 gas? A) 3.3 × 10-11 M B) 1.7 × 10-10 M C) 6.0 × 109 M D) 3.0 × 1010 M Answer: D

OpenStudy (ashley1nonly):

i suppose i have to start another question

OpenStudy (anonymous):

kk, kc is used for equilibrium and molaritys, kp is used for solubility ksp

OpenStudy (ashley1nonly):

so when we are using delta n, when both sides are equal thats when kp=kc

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kp is pressure and precipitation, kc-moles

OpenStudy (ashley1nonly):

ok

OpenStudy (anonymous):

can you pull your second problem up without the symbols

OpenStudy (ashley1nonly):

it did that by itself so i will close this one and start another question

OpenStudy (anonymous):

kk